OpenStudy (anonymous):

What Is The Equation Of The Line That Is Parallel To The Line y = 5x + 2 And Passes Through The Point (1, -3)?

OpenStudy (anonymous):

a line that is parallel to the line given has the same slope. since the slope = m, and it is in form y = mx + b (b being the y-intercept), the slope = 5. you plug this into point-slope form because you have the slope and a point of the line (1, -3). point slope form is y-y1 = m(x-x1)... y - (-3) = 5(x - 1) y+3=5x-5 y=5x-8